OpenStudy (anonymous):

I need to solve for 2x^2 -10x+3=0 using the quadratic formula. Any help?

OpenStudy (anonymous):

I am thinking, find a,b,c

OpenStudy (anonymous):

But i feel as though i am getting it wrong.

OpenStudy (welshfella):

quadratic formula is x = -b +/- sqrt (b^2 - 4ac) ------------------- 2a where a , b and c are the numbers in the equation ax^2 + bx + c = 0

OpenStudy (anonymous):

a=2 b=-10 c=3

OpenStudy (anonymous):

so plug these in ?

OpenStudy (welshfella):

right

OpenStudy (anonymous):

Ok, thanks a bunch.

OpenStudy (welshfella):

watch the signs eg -b = - (-10)

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

would -b make the -10 positive 10?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

20--24/4

OpenStudy (anonymous):

so 44 /4=10

OpenStudy (anonymous):

x=10?

OpenStudy (anonymous):

11* srry

OpenStudy (welshfella):

you left out the sqrt its (10 + sqrt76 )/ 4 or ( 10 - sqrt76) / 4 perform the calculation in the parentheses first, then divid e by 4

OpenStudy (anonymous):

OH!!!

OpenStudy (anonymous):

silly me:)

OpenStudy (anonymous):

what would be the square root i do not understand.

OpenStudy (welshfella):

use your calculator - enter square root then 76

OpenStudy (anonymous):

i see, so 11-sgrt76, but where is this 76 from?

OpenStudy (anonymous):

2.28220211292 would be the answer

OpenStudy (welshfella):

the 76 comes from the b^2 - 4ac of the formua:- 10^2 - 4*2*3 = 100 - 24 = 76

OpenStudy (anonymous):

I see. Thank you for your time. :)

OpenStudy (welshfella):

ok where did you get 2.25 from?

OpenStudy (welshfella):

2.28 and so on?

OpenStudy (anonymous):

11-sgrt76=2.28 sir

OpenStudy (welshfella):

10 - sqrt76 = 10 - 8.72 - 1.28 but you now have to divide by 4 x = 1.28 / 4 = 0.32

OpenStudy (anonymous):

hm, i see. that would be .32 OK

OpenStudy (welshfella):

thats one root and the other is (10 + sqrt76 )/ 4

OpenStudy (anonymous):

i completely forgot;/

OpenStudy (welshfella):

these quadratics can have 2 roots

OpenStudy (anonymous):

Hm.